# On the First Passage Times of Branching Random Walks in $\mathbb{R}^{d}$

### v1

@@ 2.3 Tightness of first passage times: proof of Theorem 2
On the other hand, on the event that #$\#\{v\in V_{C_{3}k}:\eta_{v,C_{3}k}(C_{3}k)\geqslant k\}\geqslant e^{c_{4}k}$, we may label $e^{c_{4}k}$ many particles by $\{v_{i}\}_{1\leqslant i\leqslant e^{c_{4}k}}$. Our goal is to show that the probability that fewer than $e^{c_{11}k}$ of $\{v_{i}\}_{1\leqslant i\leqslant e^{c_{4}k}}$ reach $B_{x}$ in the period of time $[C_{3}k,C_{3}k+s+1+C_{6}C_{7}k]$ is small. Note that the trajectories for each $i$ in the period $[C_{3}k,C_{3}k+s+1+C_{6}C_{7}k]$ are independent and equivalent in law to a BRW trajectory initiated at $\bm{\eta}_{v,C_{3}k}(C_{3}k)$. A technical difficulty arises, however, as the events of reaching $B_{x}$ are not monotone in $x$ in dimensions $d\geqslant 2$. It would be nice if we had stochastic monotonicity for $\tau_{x}$ in $x$ (as in the one-dimensional case), but this appears out of our reach. Instead, we use the following argument to bypass this technical difficulty.
@@
Recall that each $\eta_{v_{i},C_{3}k}(C_{3}k)\geqslant k$, and the notation $\mathbf{x}=(x,0\dots,0)\in\mathbb{R}^{d}$. By a standard large deviation estimate on the last $d-1$ coordinates of $\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)$ and using the triangle inequality, we may assume that $\left\lVert\mathbf{x}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)\right\rVert\leqslant x+C_{6}k$ for all $1\leqslant i\leqslant e^{c_{4}k}$ by losing an event of probability $O(e^{-c_{12}k})$ with $c_{4}$ chosen small enough. Similarly, we may assume $\eta_{v_{i},C_{3}k}(C_{3}k)\leqslant C_{8}k$ with an overwhelming probability of at least $1-O(e^{-c_{12}k})$ (since there are $e^{c_{4}k}$ many particles in total and we can apply a union bound). As a consequence,
@@
$\mathbb{P}\left(\forall 1\leqslant i\leqslant e^{c_{4}k},\leavevmode\nobreak\ x-C_{8}k\leqslant\left\lVert\mathbf{x}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)\right\rVert\leqslant x+C_{6}k\right)\geqslant 1-O(e^{-c_{12}k}).$
@@
For each $1\leqslant i\leqslant e^{c_{4}k}$, consider the set of $\mathbf{x}_{i}$’s such that $\left\lVert\mathbf{x}_{i}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)\right\rVert=x$. Selecting $C_{6}>C_{8}$ and on a set of overwhelming probability of at least $1-O(e^{-c_{12}k})$, we can select $\mathbf{x}_{i}$ such that $\left\lVert\mathbf{x}-\mathbf{x}_{i}\right\rVert\leqslant C_{8}k$. For ${\mathbf{z}}\in\mathbb{R}^{d}$, we denote by $\tau_{{\mathbf{z}}}$ the FPT to the ball of radius one centered at ${\mathbf{z}}$. By Lemma 13 applied with $C_{4}=C_{7}/2$ and picking $C_{7}$ large (and without loss of generality, a $p/2$ proportion of the $e^{c_{4}k}$ many particles do not terminate, by losing an exponentially small probability),
@@
(#$\hskip 14.22636pt\mathbb{P}\left(\#\{1\leqslant i\leqslant\frac{p}{2}e^{c_{4}k},\leavevmode\nobreak\ \tau_{\mathbf{x}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)}\geqslant s+1+C_{6}C_{7}k\}\geqslant e^{c_{11}k}\right)$
@@
(#$\leqslant\mathbb{P}\left(\#\{1\leqslant i\leqslant\frac{p}{2}e^{c_{4}k},\leavevmode\nobreak\ \tau_{\mathbf{x}_{i}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)}\geqslant s+1\}\geqslant e^{c_{11}k}\right)+O(e^{-c_{12}k})$
@@
$\hskip 28.45274pt+\mathbb{P}\left(\exists 1\leqslant i\leqslant\frac{p}{2}e^{c_{4}k},\leavevmode\nobreak\ \tau_{\mathbf{x}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)}\geqslant s+1+C_{6}C_{7}k,\leavevmode\nobreak\ \tau_{\mathbf{x}_{i}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)}<s+1\right)$
@@
(#$\leqslant\mathbb{P}\left(\#\{1\leqslant i\leqslant\frac{p}{2}e^{c_{4}k},\leavevmode\nobreak\ \tau_{\mathbf{x}_{i}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)}\geqslant s+1\}\geqslant e^{c_{11}k}\right)$
@@
Inserting back to (31) and using Lemma 13 on the second term of (31), we have
@@
For $q>0$, using the fact that (#$\mathbb{P}(\#V_{n}>0\mid S^{c})=o(e^{-c_{16}n})$ (this is a consequence of Theorem 13.3 of [5]), we obtain
$\begin{split}\mathbb{P}(\tau_{x}>s+C_{5}k)&\geqslant p\,\mathbb{P}(\tau_{x}>s+C_{5}k\mid S)+q\,\mathbb{P}(\tau_{x}>y,\leavevmode\nobreak\ \#V_{y}=0\mid S^{c})\\ &\geqslant p\,\mathbb{P}(\tau_{x}>s+C_{5}k\mid S)+q(1-o(e^{-c_{16}y}))-C_{9}\mathbb{P}(\tau_{x}\leqslant y,\leavevmode\nobreak\ S^{c}).\end{split}$ (34)
Again by using (#$\mathbb{P}(\#V_{n}>0\mid S^{c})=o(e^{-c_{16}n})$, we have for some $c_{17}$ to be determined later,
$\mathbb{P}(\tau_{x}\leqslant y,\leavevmode\nobreak\ S^{c})\leqslant o(e^{-c_{17}c_{16}y})+\mathbb{P}(\tau_{x}\leqslant c_{17}y).$
Recall that $M_{n}$ denotes the maximum in the first direction of the BRW at time step $n$. Using Lemma 5.2 of [29], it holds that for $\widetilde{M}_{j}$ denoting the maxima of #$\#V_{j}$ many independent random walks,
@@
This establishes (5). ∎
@@ 2.4 Proof of Theorem 3
First, let us recall our notations. Let $\Lambda(\bm{\lambda})=\log\phi_{\bm{\xi}}(\bm{\lambda})=\log\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]$ and $I(\mathbf{x})$ be its Fenchel-Legendre transform, which coincides with the large deviation rate function for the jump distribution ${\bm{\xi}}$. The unique positive solution to $I(\widehat{c}_{1},\mathbf{0})=\log\rho$ is given by $\widehat{c}_{1}$. Our goal is to prove that $\tau_{x}=\widehat{A}(x)+o_{\mathbb{P}}(\log x)$, where
@@
Figure 3: Visualization of how random walks in cones are related to estimating first passage times. The first passage time $\tau_{x}$ is roughly the time when the growing range (illustrated with the shaded ellipses) of the BRW becomes tangent with the target ball $B_{x}$. In our proof, we will construct moving cones (that are circular centered in the direction $-\mathbf{c}_{2}$ in the sense of (35); the boundaries are indicated by the blue lines) and show that with high probability, the range of the BRW always lies within the moving cones.
@@ 2.4.1 Random walks in cones
For estimating $\tau_{x}$, we need a few preliminary results on random walks and Brownian motion in cones. The intuitive reason why random walks in cones come into play is explained in Figure 3. In this section, we try to be self-contained and summarize only the necessary results. We refer to [6, 19, 20, 23] for more background on this topic.
@@
be the first passage time of the random walk $\{\mathbf{S}_{n}\}_{n\geqslant 1}$ to $\mathbb{K}^{\mathrm{c}}$, starting from ${\mathbf{z}}\in\mathbb{K}$. Similarly, let $\tau_{\mathbf{z}}^{\mathbb{K},\mathrm{BM}}$ be the first passage time of the standard $d$-dimensional Brownian motion to $\mathbb{K}^{\mathrm{c}}$, starting from ${\mathbf{z}}\in\mathbb{K}$. We will frequently consider cones that are circular (in the sense of [20]), that is, of the form
@@
The following result combines several results of [23].
@@
Suppose that $\mathbf{U}$ is centered, non-lattice, and $\mathbb{E}[U_{i}U_{j}]={\mathbbm{1}}_{\{i=j\}}$. Let $\mathbf{x},\mathbf{y}\in\mathbb{K}$. There exist positive functions $V,\widetilde{V}$ satisfying $\max(V(\mathbf{x}),\widetilde{V}(\mathbf{x}))\leqslant C(1+\left\lVert\mathbf{x}\right\rVert)^{p}$ and
@@
$\mathbb{P}(\mathbf{x}+\mathbf{S}_{n}\in B_{\mathbf{y}},\tau^{\mathbb{K}}_{\mathbf{x}}>n)\asymp\frac{V(\mathbf{x})}{n^{p+d/2}}.$ (37)
@@
The asymptotic (37) is a direct consequence of (1), Theorem 6, and Lemma 13(b) of [23]. On the other hand, (36) follows from the derivation of Lemmas 27 and 28 therein, along with their Theorem 1. The results in [23] are stated in the lattice case but non-lattice walks can be proved in the same way. ∎
@@
As commented in [23], if $\mathbf{U}$ does not satisfy $\mathbb{E}[U_{i}U_{j}]={\mathbbm{1}}_{\{i=j\}}$ but has a positive definite covariance matrix, then the same results hold with $\mathbb{K}$ replaced by $M\mathbb{K}=\{M\mathbf{x}:\mathbf{x}\in\mathbb{K}\}$, where $\mathbf{V}=M\mathbf{U}$ satisfies $\mathbb{E}[V_{i}V_{j}]={\mathbbm{1}}_{\{i=j\}}$.
The exponent $p(\mathbb{K})$ is in general not explicit and is related to the smallest eigenvalue of the Laplace-Beltrami operator; see [23]. However, it suffices for our purpose to understand the behavior of $p(\mathbb{K})$ when $\mathbb{K}$ is a circular cone that is close to a half-space. The following result shows that $p\approx 1$ if the circular cone $\mathbb{K}$ is close enough to a half-space.
Lemma 15 (Section 3 of [19]).
Let $\alpha\in(0,\pi)$ and ${\mathbf{z}}\in\mathbb{K}_{\alpha}$ be arbitrary. There exists a decreasing and continuous [^9] The continuity is not explicitly stated in [19], but follows from the continuity of the hypergeometric function. function $p_{\alpha,d}$ such that $p_{\pi/2,d}=1$ and
$\mathbb{E}[(\tau^{\mathbb{K}_{\alpha}}_{\mathbf{z}})^{q/2}]<\infty\quad\Longleftrightarrow\quad q\alpha<\frac{\pi}{2}.$
The following ballot-type lemma provides bounds when the cone is shifted by a function of time, borrowing ideas from Lemma 2.3 of [13]. We do not attempt the sharpest bound here, but state the minimal result needed for our purpose.
@@
Let $\mathbf{v}\in\mathbb{R}^{d},\leavevmode\nobreak\ \mathbf{v}\neq\mathbf{0}$, and $\alpha\in(0,\pi)$. Let $p=p(\mathbb{K}_{\alpha}(-\mathbf{v}))$ be the exponent defined in Lemma 14, $p^{\prime}<p$, and $C>0$. Then uniformly in ${\mathbf{z}},\mathbf{y}\in\mathbb{R}^{d}$ such that ${\mathbf{z}}-\mathbf{y}\in\mathbb{K}_{\alpha}(-\mathbf{v})$ and $n\in\mathbb{N}$,
$\begin{split}&\mathbb{P}(\mathbf{S}_{n}\in B_{\mathbf{z}};\,\forall 1\leqslant k\leqslant n,\,\mathbf{S}_{k}-\mathbf{y}-C\log(\min(k,n-k))_{+}\mathbf{v}\in\mathbb{K}_{\alpha}(-\mathbf{v}))\\ &\hskip 170.71652pt\ll\frac{(1+\left\lVert{\mathbf{z}}\right\rVert)^{p}(1+\left\lVert\mathbf{y}\right\rVert)^{p}}{n^{p^{\prime}+d/2}}.\end{split}$ (38)
Moreover, uniformly for $n\in\mathbb{N}$, $1\leqslant\ell\leqslant n$, and ${\mathbf{z}},\mathbf{y}\in\mathbb{R}^{d}$ with ${\mathbf{z}}-\mathbf{y}\in C\log(\min(\ell,n-\ell))_{+}\mathbf{v}+\mathbb{K}_{\alpha}(-\mathbf{v})$,
$\begin{split}&\mathbb{P}(\mathbf{S}_{\ell}\in B_{\mathbf{z}};\,\forall 1\leqslant k\leqslant\ell,\,\mathbf{S}_{k}-\mathbf{y}-C\log(\min(k,n-k))_{+}\mathbf{v}\in\mathbb{K}_{\alpha}(-\mathbf{v}))\\ &\hskip 170.71652pt\ll\frac{(1+\left\lVert{\mathbf{z}}\right\rVert)^{p}(1+\left\lVert\mathbf{y}\right\rVert)^{p}}{n^{p^{\prime}+d/2}}.\end{split}$ (39)

### v2

@@ 2.3 Tightness of first passage times: proof of Theorem 2
On the other hand, on the event that $\{v\in V_{C_{3}k}:\eta^{(1)}_{v,C_{3}k}(C_{3}k)\geqslant k\}\geqslant e^{c_{4}k}$, we may label $e^{c_{4}k}$ many particles by $\{v_{i}\}_{1\leqslant i\leqslant e^{c_{4}k}}$. Our goal is to show that the probability that fewer than $e^{c_{11}k}$ of $\{v_{i}\}_{1\leqslant i\leqslant e^{c_{4}k}}$ reach $B_{x}$ in the period of time $C_{3}k,C_{3}k+s+1+C_{6}C_{7}k]$ is small. Note that the trajectories for each $i$ in the period $C_{3}k,C_{3}k+s+1+C_{6}C_{7}k]$ are independent and equivalent in law to a BRW trajectory initiated at $\bm{\eta}_{v,C_{3}k}(C_{3}k)$. A technical difficulty arises, however, as the events of reaching $B_{x}$ are not monotone in $x$ in dimensions $d\geqslant 2$. It would be nice if we had stochastic monotonicity for $\tau_{x}$ in $x$ (as in the one-dimensional case), but this appears out of our reach. Instead, we use the following argument to bypass this technical difficulty.
@@
Recall that each $\eta^{(1)}_{v_{i},C_{3}k}(C_{3}k)\geqslant k$, and the notation $\mathbf{x}=(x,0\dots,0)\in\mathbb{R}^{d}$. By a standard large deviation estimate on the last d1d-1 coordinates of $\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)$ and using the triangle inequality, we may assume that $\left\lVert\mathbf{x}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)\right\rVert\leqslant x+C_{6}k$ for all $1\leqslant i\leqslant e^{c_{4}k}$ by losing an event of probability $O(e^{-c_{12}k})$ with $c_{4}$ chosen small enough. Similarly, we may assume $\eta^{(1)}_{v_{i},C_{3}k}(C_{3}k)\leqslant C_{8}k$ with an overwhelming probability $1-O(e^{-c_{12}k})$ (since there are $e^{c_{4}k}$ many particles in total and we can apply a union bound). As a consequence,
@@
$\mathbb{P}\left(\forall 1\leqslant i\leqslant e^{c_{4}k},\penalty 10000\ \max\{0,x-C_{8}k\}\leqslant\left\lVert\mathbf{x}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)\right\rVert\leqslant x+C_{6}k\right)\geqslant 1-O(e^{-c_{12}k}).$
@@
For each $1\leqslant i\leqslant e^{c_{4}k}$, consider the set of $\mathbf{x}_{i}$’s such that $\left\lVert\mathbf{x}_{i}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)\right\rVert=x$. Selecting $C_{8}>C_{6}$ and on a set of overwhelming probability of at least $1-O(e^{-c_{12}k})$, we can select $\mathbf{x}_{i}$ such that $\left\lVert\mathbf{x}-\mathbf{x}_{i}\right\rVert\leqslant C_{8}k$. For ${\mathbf{z}}\in\mathbb{R}^{d}$, we denote by $\tau_{{\mathbf{z}}}$ the FPT to the ball of radius one centered at $\mathbf{z}$}. By Lemma 13 applied with $C_{4}=C_{7}/2$ and picking $C_{7}$ large (and without loss of generality, a $p/2$ proportion of the $e^{c_{4}k}$ many particles do not terminate, by losing an exponentially small probability),
@@
$\hskip 14.22636pt\mathbb{P}\left(\#\{1\leqslant i\leqslant\frac{p}{2}e^{c_{4}k},\penalty 10000\ \tau_{\mathbf{x}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)}\geqslant s+1+C_{6}C_{7}k\}\geqslant e^{c_{11}k}\right)$
@@
$\leqslant\mathbb{P}\left(\#\{1\leqslant i\leqslant\frac{p}{2}e^{c_{4}k},\penalty 10000\ \tau_{\mathbf{x}_{i}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)}\geqslant s+1\}\geqslant e^{c_{11}k}\right)+O(e^{-c_{12}k})$
@@
$\hskip 28.45274pt+\mathbb{P}\left(\exists 1\leqslant i\leqslant\frac{p}{2}e^{c_{4}k},\penalty 10000\ \tau_{\mathbf{x}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)}\geqslant s+1+C_{6}C_{7}k,\penalty 10000\ \tau_{\mathbf{x}_{i}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)}<s+1\right)$
@@
$\leqslant\mathbb{P}\left(\#\{1\leqslant i\leqslant\frac{p}{2}e^{c_{4}k},\penalty 10000\ \tau_{\mathbf{x}_{i}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)}\geqslant s+1\}\geqslant e^{c_{11}k}\right)$
@@
Inserting back to (31) and using Lemma 13 on the third term on the right-hand side of (31) (with our choice of $C_{5}\gg 1+C_{3}+C_{6}C_{7}$), we have
@@
For $q>0$, using the fact that $\mathbb{P}(\#V_{n}>0\mid S^{c})=o(e^{-c_{16}n})$ (this is a consequence of Theorem 13.3 of [athreya2004branching]), we obtain
$\begin{split}\mathbb{P}(\tau_{x}>s+C_{5}k)&\geqslant p\,\mathbb{P}(\tau_{x}>s+C_{5}k\mid S)+q\,\mathbb{P}(\tau_{x}>y,\penalty 10000\ \#V_{y}=0\mid S^{c})\\ &\geqslant p\,\mathbb{P}(\tau_{x}>s+C_{5}k\mid S)+q(1-o(e^{-c_{16}y}))-C_{9}\mathbb{P}(\tau_{x}\leqslant y,\penalty 10000\ S^{c}).\end{split}$ (34)
Again by using $\mathbb{P}(\#V_{n}>0\mid S^{c})=o(e^{-c_{16}n})$, we have for some $c_{17}$ to be determined later,
$\mathbb{P}(\tau_{x}\leqslant y,\penalty 10000\ S^{c})\leqslant o(e^{-c_{17}c_{16}y})+\mathbb{P}(\tau_{x}\leqslant c_{17}y).$
Recall that $M_{n}$ denotes the maximum in the first direction of the BRW at time step $n$. Using Lemma 5.2 of [gantert2018large], it holds that for $\widetilde{M}_{j}$ denoting the maxima of $V_{j}$ many independent random walks,
@@
For the case $y\leqslant 2C_{5}y_{0}/C_{3}$, we may increase the constant $C$ in (5) such that for $y$ in this range, $Ce^{-cy}\geqslant 1$ holds, and hence (5) trivially holds. This establishes (5). ∎
We remark that the above proof works for a general $r\in(0,1)$ (where $r$ is the quantile in $t_{x}^{(r)}$), in addition to the case $r=1/2$.
@@ 2.4 Proof of Theorem 3
First, let us recall our notations. Let $\Lambda(\bm{\lambda})=\log\phi_{\bm{\xi}}(\bm{\lambda})=\log\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]$ and $I(\mathbf{x})$ be its Fenchel–Legendre transform, which coincides with the large deviation rate function for the jump distribution $\bm{\xi}$}. The unique positive solution to $I(\widehat{c}_{1},\mathbf{0})=\log\rho$ is given by $\widehat{c}_{1}$. Our goal is to prove that $\tau_{x}=\widehat{A}(x)+o_{\mathbb{P}}(\log x)$, where
@@
Figure 3: Visualization of how random walks in cones are related to estimating first passage times. The first passage time $\tau_{x}$ is roughly the time when the growing range (illustrated with the shaded ellipses) of the BRW becomes tangent with the target ball $B_{x}$. In our proof, we will construct moving cones (that are circular centered in the direction $\mathbf{c}_{2}$ in the sense of (35); the boundaries are indicated by the blue lines) and show that with high probability, the range of the BRW always lies within the moving cones.
@@ 2.4.1 Random walks in cones
For estimating $\tau_{x}$, we need a few preliminary results on random walks and Brownian motion in cones. The intuitive reason why random walks in cones come into play is explained in Figure 3. In this section, we try to be self-contained and summarize only the necessary results. We refer to [banuelos1997brownian, burkholder1977exit, deblassie1987exit, denisov2015random] for more background on this topic.
@@
be the first passage time of the random walk $\{\mathbf{S}_{n}\}_{n\geqslant 1}$ to $\mathbb{K}^{\mathrm{c}}$, starting from ${\mathbf{z}}\in\mathbb{K}$. Similarly, let $\tau_{\mathbf{z}}^{\mathbb{K},\mathrm{BM}}$ be the first passage time of the standard $d$-dimensional Brownian motion to $\mathbb{K}^{\mathrm{c}}$, starting from ${\mathbf{z}}\in\mathbb{K}$. We will frequently consider cones that are circular (in the sense of [deblassie1987exit]), that is, of the form
@@
The following result combines several results of [denisov2015random]. Recall from assumption (A5) that the law of $\bm{\xi}$} is non-lattice.
@@
Fix a circular cone $\mathbb{K}$, i.e., $\mathbb{K}=\mathbb{K}_{\alpha}$ for some $\alpha\in(0,\pi)$. Suppose that $\mathbf{U}$ is centered, non-lattice, and $\mathbb{E}[U_{i}U_{j}]={\mathbbm{1}}_{\{i=j\}}$. Let $\mathbf{x},\mathbf{y}\in\mathbb{K}$. There exist positive functions $V,\widetilde{V}$ satisfying $\max(V(\mathbf{x}),\widetilde{V}(\mathbf{x}))\leqslant C(1+\left\lVert\mathbf{x}\right\rVert)^{p}$ and
@@
$\mathbb{P}(\mathbf{x}+\mathbf{S}_{n}\in B_{\mathbf{y}},\tau^{\mathbb{K}}_{\mathbf{x}}>n)\asymp\frac{V(\mathbf{x})}{n^{p+d/2}},$ (37)
where the asymptotic constant in $\asymp$ may depend on $y$.
@@
The asymptotic (37) is a direct consequence of (1), (10) of Theorem 6 (which holds also for a fixed $\mathbf{y}\in\mathbb{K}$), and Lemma 13(b) of [denisov2015random]. On the other hand, (36) follows from the derivation of Lemmas 27 and 28 therein, along with their Theorem 1. The results in [denisov2015random] are stated in the lattice case but non-lattice walks can be proved in the same way. ∎
@@
As commented in [denisov2015random], if $\mathbf{U}$ does not satisfy $\mathbb{E}[U_{i}U_{j}]={\mathbbm{1}}_{\{i=j\}}$ but has a positive definite covariance matrix, then the same results hold with $\mathbb{K}$ replaced by $M\mathbb{K}=\{M\mathbf{x}:\mathbf{x}\in\mathbb{K}\}$, where $\mathbf{V}=M\mathbf{U}$ satisfies $\mathbb{E}[V_{i}V_{j}]={\mathbbm{1}}_{\{i=j\}}$.
The exponent $p(\mathbb{K})$ is in general not explicit and is related to the smallest eigenvalue of the Laplace-Beltrami operator; see [denisov2015random]. However, it suffices for our purpose to understand the behavior of $p(\mathbb{K})$ when $\mathbb{K}$ is a circular cone that is close to a half-space. The following result shows that $p\approx 1$ if the circular cone $\mathbb{K}$ is close enough to a half-space.
Lemma 15 ((3.11) of [burkholder1977exit]).
Let $\alpha\in(0,\pi)$ and ${\mathbf{z}}\in\mathbb{K}_{\alpha}$ be arbitrary. There exists a decreasing and continuous function $p_{\alpha,d}$ such that $p_{\pi/2,d}=1$ and [^10] The formal definition is given in [burkholder1977exit] by $p_{\alpha,d}=\sup\{p:\alpha<\theta_{p,d}\}$, where $\theta_{p,d}$ is the smallest zero in $(0,\pi)$ of $h(\theta)=F(-p,p+d-2;(d-1)/2;(1-\cos\theta)/2)$. The continuity is not explicitly stated in [burkholder1977exit], but follows from the continuity of the hypergeometric function.
$\mathbb{E}[(\tau^{\mathbb{K}_{\alpha}}_{\mathbf{z}})^{q/2}]<\infty\quad\Longleftrightarrow\quad q<p_{\alpha,d}.$
The following ballot-type lemma provides bounds when the cone is shifted by a function of time. We do not attempt the sharpest bound here, but state the minimal result needed for our purpose.
@@
Assume the same setting on the law of $\mathbf{U}$ as in Lemma 14. Let $\mathbf{v}\in\mathbb{R}^{d},\penalty 10000\ \mathbf{v}\neq\mathbf{0}$, and $\alpha\in(0,\pi)$. Let $p=p(\mathbb{K}_{\alpha}(-\mathbf{v}))$ be the exponent defined in Lemma 14, $p^{\prime}<p$, and $C>0$. Then uniformly in ${\mathbf{z}},\mathbf{y}\in\mathbb{R}^{d}$ such that ${\mathbf{z}}-\mathbf{y}\in\mathbb{K}_{\alpha}(-\mathbf{v})$ and $n\in\mathbb{N}$,
$\begin{split}&\mathbb{P}(\mathbf{S}_{n}\in B_{\mathbf{z}};\,\forall 1\leqslant k\leqslant n,\,\mathbf{S}_{k}-\mathbf{y}-C\log(\min(k,n-k))_{+}\mathbf{v}\in\mathbb{K}_{\alpha}(-\mathbf{v}))\\ &\hskip 170.71652pt\ll\frac{(1+\left\lVert{\mathbf{z}}\right\rVert)^{p}(1+\left\lVert\mathbf{y}\right\rVert)^{p}}{n^{p^{\prime}+d/2}},\end{split}$ (38)
where the asymptotic constant in $\ll$ depends on $d,p^{\prime}$, and the law of $\bm{\xi}$}. Moreover, uniformly for $n\in\mathbb{N}$, $1\leqslant\ell\leqslant n$, and ${\mathbf{z}},\mathbf{y}\in\mathbb{R}^{d}$ with ${\mathbf{z}}-\mathbf{y}\in C\log(\min(\ell,n-\ell))_{+}\mathbf{v}+\mathbb{K}_{\alpha}(-\mathbf{v})$,
$\begin{split}&\mathbb{P}(\mathbf{S}_{\ell}\in B_{\mathbf{z}};\,\forall 1\leqslant k\leqslant\ell,\,\mathbf{S}_{k}-\mathbf{y}-C\log(\min(k,n-k))_{+}\mathbf{v}\in\mathbb{K}_{\alpha}(-\mathbf{v}))\\ &\hskip 170.71652pt\ll\frac{(1+\left\lVert{\mathbf{z}}\right\rVert)^{p}(1+\left\lVert\mathbf{y}\right\rVert)^{p}}{n^{p^{\prime}+d/2}},\end{split}$ (39)
where the asymptotic constant in $\ll$ depends on $d,p^{\prime}$, and the law of $\bm{\xi}$}.